# Time, Travel, and Energy in the Uniform Dispersion Problem

### v1

@@ Abstract
We investigate the algorithmic problem of uniformly dispersing a swarm of robots in an unknown, gridlike environment. In this setting, our goal is to comprehensively study the relationships between performance metrics and robot capabilities. We introduce a formal model comparing dispersion algorithms based on makespan, traveled distance, energy consumption, sensing, communication, and memory. Using this framework, we classify several uniform dispersion algorithms according to their capability requirements and performance. We prove that while makespan and travel can be minimized in all environments, energy cannot, as long as the swarm’s sensing range is bounded. In contrast, we show that energy can be minimized even by simple, “ant-like” robots in synchronous settings and asymptotically minimized in asynchronous settings, provided the environment is topologically simply connected. Our findings offer insights into fundamental limitations that arise when designing swarm robotics systems for exploring unknown environments, highlighting the impact of environment’s topology on the feasibility of energy-efficient dispersion.
@@ I Introduction
A notable approach to using swarms for exploration and monitoring is the concept of ”flooding” or ”uniform dispersion” [3, 4, 5, 6, 7, 8, 9, 10]. This approach involves gradually deploying a swarm of small, expendable robotic agents at a fixed entry point in the environment and allowing them to explore and settle until the area is fully covered. By operating autonomously and in parallel, these robots can efficiently gather crucial data while reducing the risk to human personnel [8]. Potential applications range from search and rescue missions and environmental monitoring to scientific exploration.
In this work, we study the algorithmic problem the heart of such an approach: flooding an unknown, discrete grid-like environment with a swarm of robots that gradually enter from a source location. This problem, known in the literature as the uniform dispersion problem, was formally posed by Hsiang et al. [3] (building on the earlier work of Howard et al. [11]) and has attracted significant attention in the literature–see “Related Work”. By studying the uniform dispersion problem in a formal mathematical setting, we can gain insights into the fundamental trade-offs and limitations of swarm robotics algorithms, which can inform the design of such systems.
@@
Since cheap and expendable robots are central to swarm robotics, in addition to these performance metrics, a good uniform dispersion algorithm should strive to minimize the hardware requirements for successful execution, such as sensing range, communication bandwidth (using local signals), and memory or computation capabilities [12, 13, 14, 15]. Optimizing these metrics and minimizing the required capabilities raises important questions about the relationships between them:
@@
This work aims to be a comprehensive investigation of such questions in a uniform dispersion setting. We introduce a formal model that enables us to compare uniform dispersion algorithms in terms of three performance metrics: makespan, travel, and energy, and three capability requirements: sensing, communication, and persistent state memory (a proxy for the dispersion strategy’s complexity). Robots with sensing range $V$, communication bandwidth $B$ and persistent state memory $S$ are described as $(V,B,S)$-robots. Using this framework, we identify sufficient robot capabilities for minimizing each performance metric. The results are summarized in Table I.
TABLE I: A summary of uniform dispersion results by performance metric and environment type.
@@
Energy General Cannot minimize assuming constant $V$
Energy Simply Connected (2, 0, 5)-robots
(2, 1, 5)-robots (Asynchronously)
Our results show that, although there exist uniform dispersion algorithms that minimize makespan (a result due to Hsiang et al. [3]) and travel (our own result) in general environments, there does not exist an algorithm that minimizes total energy in general environments, assuming robots do not know the environment in advance (in other words, $(V,\infty,\infty)$-robots cannot minimize energy for any given, constant $V$). In fact, we prove that even a centralized algorithm with unlimited computational power but without prior knowledge of the environment cannot minimize energy (Section III). Informally, this is because an energy-optimal algorithm necessarily minimizes both makespan and travel, hence energy necessitates both deciding quickly and optimal path-finding–two goals in conflict with each other.
Given the above, we are led to ask whether energy can be minimized in more restricted types of environments. We show that a sufficient topological condition is simply-connectedness [16]. A simply connected environment is one where any closed loop can be continuously shrunk to a point without leaving the environment. Informally, this means the environment has no “holes” or “obstacles” that a loop could get stuck around. Examples of simply connected environments include convex shapes like rectangles, as well as more complex shapes like mazes or building floors, as long as they don’t have any enclosed holes. We describe an algorithm, “Find-Corner Depth-First Search” (FCDFS), that minimizes energy in such environments. FCDFS is an algorithm for $(2,0,5)$-robots, meaning it requires no communication and has small, constant sensing range and memory requirements. This makes it executable by very simple “ant-like” robots, in contrast with our impossibility results in non-simply connected environments.
@@
The results we described above all assume a synchronous setting where robots all activate simultaneously at every time step, but in real world applications, we expect asynchronicity. It is thus important to ask whether FCDFS generalizes to asynchronous settings. We show that it does in Section V, where we consider an asynchronous setting with random robot activation times. We show that, asymptotically, an asynchronous variant of FCDFS remains time, travel, and energy-efficient in this setting. To prove this, we relate FCDFS to the “Totally Asymmetric Simple Exclusion Process” (TASEP) in statistical physics, studied extensively in e.g. [22, 23, 24]. This proof technique originally appeared in [7] in the context of makespan. We adapt the ideas in [7] and extend them to also bound (maximal individual) energy.
@@ I-A Related Work
A large body of work exists on deploying robotic swarms for covering, exploring, or monitoring unknown environments [2]. Decentralized approaches are often favored, as they tend to be more adaptive than centralized control to uncertain environmental conditions, e.g., [25, 26, 27, 11, 3, 10, 4, 8, 28]. Among these approaches, significant attention has been given to uniform dispersion in particular. Foundational work on uniform dispersion can be traced back to Hsiang et al. [3] and Howard et al. [11]. Substantial subsequent developments focused on provably guaranteeing uniform dispersion under conditions such as asynchronicity, reduced or corrupted memory, myopic sensing capabilities, and adaptation to sudden robot crashes [29, 5, 30, 31, 6, 7, 4, 8]. Works on dispersion and similar problems have either considered environments as continuous [20, 11, 26, 32] or discrete environments [29, 33], with our work leaning towards the latter for its simplicity in representing complex environments, and due to our focusing on global performance efficiencies rather than local dynamics.
Energy is a primary constraint in robotics, and consequently, many works have dealt with predicting, or optimizing energy use in swarms [34, 21]. Whereas Hsiang et al. focused on makespan and travelled distance as optimization targets [3], the first work to consider activity time as a proxy for energy use in a uniform dispersion context is [15] (whose results this present work builds upon). Subsequently, in [8, 35], redundancy - i.e., having more robots than necessary - is shown to sometimes enable energy savings in a dual-layer uniform dispersion strategy.
@@
To bound the makespan and energy use of an asynchronous version of our energy-optimal algorithm, we relate it to the “Totally Asymmetric Simple Exclusion Process” (TASEP) in statistical physics, studied extensively in e.g. [22, 23, 24]. This process has been used to study traffic flow [37] and biological transport [38]. We are particularly interested in what is known as TASEP with step initial conditions. Two accessible references on this that have been particularly helpful to us are Romik, 2015 [39] and Kriecherbauer and Krug, 2020 [40].
@@ II Model
In the uniform dispersion problem, a swarm of autonomous robots (also called “agents”) $A_{1},A_{2},\ldots$ are tasked with completely filling an a priori unknown discrete region $R$, attempting to reach a state such that every location in $R$ contains a robot. We assume $R$ is a subset of size $n$ of the infinite grid $\mathbb{Z}^{2}=\mathbb{Z}\times\mathbb{Z}$, and represent its locations via coordinates $(x,y)$ where $x$ and $y$ are both integers. Two locations $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are connected if and only if the Manhattan distance $|x_{1}-x_{2}|+|y_{1}-y_{2}|$ is exactly $1$. The complement of $R$, denoted $R^{c}$, is defined as the subregion $\mathbb{Z}^{2}-R$. We call the vertices of $R^{c}$ walls.
Each robot (or “agent”) is represented as a mobile point in $R$. Robots are dispersed over time onto $R$ and move within it. Robots are identical and anonymous, are initialized with identical orientation, and use the same local algorithm to decide their next movements. Collisions are disallowed: a robot may not move to a location occupied by another robot, nor may two robots move to the same location simultaneously.
@@
All robots are initially active, and eventually settles. Settled robots never move from their current position. In other words, a robot “settles” after it arrives at its final, desired location. A robot may not move and settle in the same time step.
We wish to study warms of autonomous robots whose goal is to attain uniform dispersion while minimizing three performance metrics: makespan, travel, and energy.
@@
Whereas travel measures the distance a robot has moved since its arrival, energy measures the amount of time during which it was active and consumed energy. Hence, a robot’s energy use continues increasing even when when it stays put, as long as it isn’t settled:
@@
We are interested in studying autonomous robots with limited visibility that do not know their environment in advance. Primarily, we are interested in simple robots that have highly limited sensing, computation, and local signaling capabilities (as well as robots that cannot communicate at all). By “local signalling” we mean that each robot can broadcast a local visual or auditory signal by physical means, for example by emitting lights or sounds, which is received by all robots that sense it, and those robots can identify the robot that sent the message. This communication scheme is distinct from, e.g., radio communication. Let us parametrize the robots’ capabilities as follows:
@@
    Visibility range ($V$). A robot located at $p=(x,y)$ senses locations of $\mathbb{Z}^{2}$ at Manhattan distance $V$ or less from $p$. The robot knows the position relative to $p$ of every vertex of $p^{\prime}\in\mathbb{Z}^{2}$ it senses (i.e., it knows $p^{\prime}-p$). We assume the robot can tell whether $p^{\prime}$ is $s$ and whether $p^{\prime}$ contains an obstacle (i.e., another robot or a wall). However, it cannot tell from sensing alone whether an obstacle is a wall, a settled robot, or an active robot.
@@
    Communication bandwidth ($B$). Each robot can broadcast a binary string of length $B$ to all robots that sense it. When $B=0$, robots cannot broadcast.
@@
    Persistent state memory ($S$). Each robot has a state memory of $S$ bits that persists from time step to time step (meaning at a given time $t$ each robot can be in any of $2^{S}$ states).
@@
With the above capability measures in mind, we can characterize a robot’s capabilities by a 3-tuple $(V,B,S)$. A “powerful” robot might have very large $V$, $B$, or $S$, whereas a simple robot has small $V$, $B$, and $S$. The parameters $V$, $B$ and $S$ determine the possible algorithms robots can run. A central goal of ours is to study what algorithms become available to the designer of the swarm as we increase the values $V$, $B$ or $S$. When can increasing these parameters improve the quality of the robots’ dispersion strategy, by enabling them to minimize a given performance metric?

### v2

@@ Abstract
We investigate the algorithmic problem of uniformly dispersing a swarm of robots in an unknown, gridlike environment. In this setting, our goal is to study the relationships between performance metrics and robot capabilities. We introduce a formal model comparing dispersion algorithms based on makespan, traveled distance, energy consumption, sensing, communication, and memory. Using this framework, we classify uniform dispersion algorithms according to their capability requirements and performance. We prove that while makespan and travel can be minimized in all environments, energy cannot, if the swarm’s sensing range is bounded. In contrast, we show that energy can be minimized by “ant-like” robots in synchronous settings and asymptotically minimized in asynchronous settings, provided the environment is topologically simply connected, by using our “Find-Corner Depth-First Search” (FCDFS) algorithm. Our theoretical and experimental results show that FCDFS significantly outperforms known algorithms. Our findings reveal key limitations in designing swarm robotics systems for unknown environments, emphasizing the role of topology in energy-efficient dispersion.
Index Terms:
Swarm robotics, energy use, autonomous agents, co-design, formal analysis
@@ I Introduction
A notable approach to using swarms for exploration and monitoring is the concept of “flooding” or “uniform dispersion” [3, 4, 5, 6, 7, 8, 9, 10]. This approach involves gradually deploying a swarm of small, expendable robotic agents at a fixed entry point in the environment and allowing them to explore and settle until the area is fully covered. By operating autonomously and in parallel, these robots can efficiently gather crucial data while reducing the risk to human personnel [8]. Potential applications range from search and rescue missions and environmental monitoring to scientific exploration.
In this work, we study the algorithmic problem at the heart of such an approach: flooding an unknown, discrete grid-like environment with a swarm of robots that gradually enter from a source location. This problem, known in the literature as the uniform dispersion problem, was formally posed by Hsiang et al. [3] (building on the earlier work of Howard et al. [11]) and has attracted significant attention in the literature–see “Related Work”. Effective uniform dispersion algorithms are the critical algorithmic backbone of robotic flooding.
Performance Metrics and Robot Capabilities
@@
Since cheap and expendable robots are central to swarm robotics, in addition to these performance metrics, a good uniform dispersion algorithm should strive to minimize the hardware requirements for successful execution, such as sensing range, communication bandwidth (using local signals), and memory or computation capabilities [12, 13, 14, 15].
Existing Solutions and Their Limitations
Prior work on uniform dispersion has made significant progress in several areas:
  1. 1.
    Makespan optimization: Hsiang et al. [3] demonstrated algorithms that achieve optimal makespan using simple robots with constant sensing range and memory.
  2. 2.
    Robustness: Researchers have developed algorithms that guarantee uniform dispersion under challenging conditions like asynchronicity, reduced memory, myopic sensing, and sudden robot crashes [16, 5, 17, 18, 6, 7, 4, 8].
  3. 3.
    Different settings and robot dynamics: Various approaches have been proposed for both continuous [19, 11, 20, 21] and discrete environments [16, 8], with varying assumptions about the robots.
However, existing solutions have several key limitations. Most prior work has focused on optimizing individual metrics (typically makespan) without considering multiple performance measures simultaneously. Moreover, the relationship between robot capabilities and achievable performance has not been systematically studied across many metrics. Energy efficiency, while crucial for practical applications [22, 23], has received particularly limited attention. Additionally, the impact of environment topology on achievable performance remains poorly understood, leaving open questions about how environmental constraints affect optimization possibilities.
These limitations point to a significant research gap: the lack of a comprehensive theoretical framework for understanding how robot capabilities affect their ability to optimize different performance metrics in uniform dispersion. Specifically:
  • We lack formal methods for comparing uniform dispersion algorithms based on both their capability requirements and performance characteristics.
  • The fundamental relationships between makespan, travel, and energy optimization remain unclear.
  • The minimal robot capabilities needed to achieve different optimization goals have not been systematically determined.
  • The role of environment topology in enabling or constraining performance optimization is not well understood.
This gap has practical implications for swarm robotics system design, as it leaves open questions about what level of robot sophistication is truly necessary for different performance objectives.
Research Questions and Approach
To address these gaps, we investigate the following key questions:
@@
We analyze these questions from a formal, mathematical lens. We introduce a capability-based model for comparing uniform dispersion algorithms and use it to prove fundamental results about which performance metrics can be optimized by robots with different capability levels. This theoretical framework reveals fundamental relationships between robot capabilities, environment topology, and achievable performance, with notable implications for swarm robotics system design.
Specifically, we compare uniform dispersion algorithms in terms of three performance metrics: makespan, travel, and energy, and three capability requirements: sensing, communication, and persistent state memory (a proxy for the dispersion strategy’s complexity). Robots with sensing range $V$ (how many grid cells they can see around them), communication bandwidth $B$ (how many bits they can broadcast per time step), and persistent state memory $S$ (how many bits of memory persist between time steps) are described as $(V,B,S)$-robots. For example, a $(2,0,5)$-robot can see two cells in each direction, cannot communicate, and maintains 5 bits of state memory. With this framework, we identify sufficient robot capabilities for attaining optimal performance in each performance metric. Full details of the robot capability model are provided in Section II. The results are summarized in Table I.
TABLE I: Summary of uniform dispersion results by performance metric and environment type.
@@
Energy General Not possible assuming constant $V$
All metrics Simply Connected (2, 0, 5)-robots (Synchronously)
(simultaneously) (2, 1, 5)-robots (Asynchronously)
Our results show that, although there exist uniform dispersion algorithms that minimize makespan (a result due to Hsiang et al. [3]) and travel (our own result) in general environments, there does not exist an algorithm that minimizes total energy in general environments, assuming robots do not know the environment in advance (in other words, $(V,\infty,\infty)$-robots cannot minimize energy for any given, constant $V$). In fact, we prove that even a centralized algorithm with unlimited computational power but without prior knowledge of the environment cannot minimize energy (Section III). Informally, this is because an energy-optimal algorithm simultaneously also minimizes both makespan and travel, hence energy necessitates both deciding quickly and optimal path-finding–two goals in conflict with each other.
Given this conclusion, we are led to ask whether energy can be minimized in more restricted types of environments. We show that a sufficient topological condition is simply-connectedness [24]. A simply connected environment is one where any closed loop can be continuously shrunk to a point without leaving the environment, i.e., the environment has no enclosed “holes” that a loop could get stuck around. Examples of simply connected environments include convex shapes like rectangles, as well as more complex shapes like mazes or building floors, as long as they don’t have any enclosed holes. We describe an algorithm, “Find-Corner Depth-First Search” (FCDFS), that minimizes energy–thus also makespan and travel–in such environments. The idea of FCDFS is to maintain a geometric invariant: the environment must stay simply connected whenever a robot settles. FCDFS is an algorithm for $(2,0,5)$-robots, meaning it requires no communication and has small, constant sensing range and memory requirements. The $5$ bits of memory are used to store information about the previous steps of the agent in a highly compressed way, enabling it to exploit aspects of simply connected topologies. FCDFS is presented and studied in Section IV (Algorithm 1). Specifically, Section IV-B shows it can be implemented with $5$ bits of memory. Such low requirements make it executable by very simple “ant-like” robots, in contrast with our impossibility results in non-simply connected environments.
@@
The results we described above all assume a synchronous setting where robots all activate simultaneously at every time step, but in real world applications, we expect asynchronicity. It is thus important to ask whether FCDFS generalizes to asynchronous settings. We show that it does in Section V, where we consider an asynchronous setting with random robot activation times, relaxing many of our structural assumptions regarding, e.g., inter-robot distance. We show that, asymptotically, an asynchronous variant of FCDFS remains time, travel, and energy-efficient. To prove this, we relate FCDFS to the “Totally Asymmetric Simple Exclusion Process” (TASEP) in statistical physics, studied extensively in e.g. [28, 29, 30].
Summary of Contributions
  1. 1.
    A formal model for comparing uniform dispersion algorithms based on robot capabilities (sensing range, communication bandwidth, persistent memory), which guides swarm robotics design by identifying the minimal capabilities needed to achieve different performance objectives.
  2. 2.
    Fundamental impossibility results showing that while makespan and travel can be minimized individually in all environments by simple robots, energy (which requires minimizing them simultaneously) cannot be minimized even by a centralized algorithm with unlimited computational power. This reveals inherent limitations in swarm robotics design: in general environments, perfect energy efficiency is unattainable regardless of robot capabilities.
  3. 3.
    The FCDFS algorithm, which provably achieves energy-optimal dispersion in simply connected environments using minimal robot capabilities (no communication, small sensing range, and only a few bits of memory). This demonstrates that in certain common environments like convex shapes and mazes, energy-efficient dispersion is achievable even with very simple, cheap robots.
  4. 4.
    AsynchFCDFS, an asynchronous variant of FCDFS that provably maintains energy efficiency under timing conditions where robots activate randomly and asynchronously. We provide rigorous performance guarantees that hold even with unreliable robot activation.
  5. 5.
    Theoretical results characterizing the distinction between centralized algorithms, decentralized algorithms, and “strongly” decentralized ant-like algorithms. These results inform the design choice between sophisticated centralized control and simple decentralized strategies.
  6. 6.
    In addition to the theoretical performance guarantees for FCDFS and AsynchFCDFS, we present comprehensive empirical evaluation (Section VI) demonstrating that FCDFS significantly outperforms existing algorithms in the literature (BFLF and DFLF [3]), while AsynchFCDFS maintains performance within constant factors of optimal even with random robot activation. This evaluation affirms our theoretical results, further validating our claims.
Our results emphasize the importance of co-design in swarm robotics: rather than building increasingly sophisticated robots, designers should carefully match robot capabilities to environmental constraints. We show that in many practical scenarios, very simple robots can achieve optimal performance if the environment permits it.
Furthermore, our results address a question first raised by Hsiang et al. [3] about simultaneously minimizing makespan and robot travel. When considering energy use (which includes time spent stationary while active), we prove such simultaneous optimization is impossible in general environments, but achievable in simply connected environments–even by simple ’ant-like’ robots using only $5$ bits of memory and no communication. The case where time spent stationary is not counted remains open.
@@ I-A Related Work
A large body of work exists on deploying robotic swarms for covering, exploring, or monitoring unknown environments [2]. Decentralized approaches are often favored, as they tend to be more adaptive than centralized control to uncertain environmental conditions, e.g., [31, 20, 32, 11, 3, 10, 4, 8, 33]. Among these approaches, significant attention has been given to uniform dispersion in particular. Foundational work on uniform dispersion can be traced back to Hsiang et al. [3] and Howard et al. [11]. Substantial subsequent developments focused on provably guaranteeing uniform dispersion under conditions such as asynchronicity, reduced or corrupted memory, myopic sensing capabilities, and adaptation to sudden robot crashes [16, 5, 17, 18, 6, 7, 4, 8]. Works on dispersion and similar problems have either considered environments as continuous [19, 11, 20, 21] or discrete environments [16, 34], with our work choosing the latter. As our goal in this work is to characterize the algorithmic aspects of makespan, travel, and energy efficiency in uniform dispersion, we use discrete grid-like environments that abstract away continuous motion dynamics and geometric nuances. This type of abstraction is useful and common in many robotic systems, such as warehouse robotics [35].
Energy is a primary constraint in robotics, and consequently, many works have dealt with predicting, or optimizing energy use in swarms [22, 23]. Whereas Hsiang et al. focused on makespan and travelled distance as optimization targets [3], the first work to consider activity time as a proxy for energy use in a uniform dispersion context is [15] (whose results this present work builds upon). Subsequently, in [8], redundancy (having more robots than necessary) is shown to sometimes enable energy savings in a dual-layer uniform dispersion strategy.
Recent robotics’ work on travel and energy use has focused on planning and navigation in diverse contexts, including adaptive path planning for environment monitoring [36], formation-based cooperative reconnaissance [37], integrated path planning with speed control [38], and multi-objective scheduling [39] and drone-routing [40]. Unlike our setting, many of these methods rely on full knowledge of the map, and develop empirical strategies for particular scenarios. Our work focuses, instead, on theoretical foundations: we establish theoretical bounds on energy optimization in uniform dispersion, proving it impossible in unknown environments regardless of robot capabilities. This impossibility result, alongside our formal framework relating robot capabilities to achievable performance, is complementary to such approaches.
@@
To bound the makespan and energy use of AsynchFCDFS, we connect it to the “Totally Asymmetric Simple Exclusion Process” (TASEP), a particle process in statistical physics [28, 29, 30]. While TASEP has been applied to several domains, including traffic flow [42] and biological transport [43], to our knowledge–its only prior appearance in swarm robotics is our earlier conference paper [7], where a simpler single-active-agent model was analyzed and the coupling served solely to obtain a high-probability makespan bound. The present work advances that idea in two fundamental ways. First, our asynchronous setting allows multiple robots to move concurrently, and we must account for that in our analysis. Second–and most importantly–we show how we can use the connection to TASEP to bound the cumulative activity time of every robot, yielding provable energy guarantees (Proposition V.2). Readers interested in the mathematical foundations of TASEP are recommended the introductory works [44] and [45].
@@ II Model
In the uniform dispersion problem, a swarm of autonomous robots (also called “agents”) $A_{1},A_{2},\ldots$ are tasked with completely filling an a priori unknown discrete region $R$, attempting to reach a state such that every location in $R$ contains a robot. We assume $R$ is a subset of size $n$ of the infinite grid $\mathbb{Z}^{2}=\mathbb{Z}\times\mathbb{Z}$, and represent its locations via coordinates $(x,y)$ where $x$ and $y$ are both integers. Two locations $(x_{1},y_{1})$ and $(x_{2},y_{2})$ are connected if and only if the Manhattan distance $|x_{1}-x_{2}|+|y_{1}-y_{2}|$ is exactly $1$ (so diagonal locations are not connected). The complement of $R$, denoted $R^{c}$, is defined as the subregion $\mathbb{Z}^{2}-R$. We call the vertices of $R^{c}$ walls.
Each robot (or “agent”) is represented as a mobile point in $R$. Robots are dispersed over time onto $R$ and move within it between connected locations, always occupying a single grid-cell of $R$. Robots are identical and anonymous, are initialized with identical orientation, maintaining a global sense of orientation through dead-reckoning or a compass, and use the same local algorithm to decide their next movements. A robot should not move to a location already occupied by another robot, nor should two robots move to the same location simultaneously. We do not assume such collision avoidance, rather, the algorithms we study are designed and formally proven to satisfy this requirement.
Note that in this discrete grid model, both robots and walls simply occupy integer coordinates in $\mathbb{Z}^{2}$, with no notion of continuous shapes or sizes. This discrete representation simplifies both sensing and movement: robots need only detect whether neighboring grid cells are occupied or unoccupied, and movement consists of transitioning between adjacent grid cells.
@@
All robots are initially active, and eventually settle. Settled robots never move from their current position. In other words, a robot “settles” after it arrives at its final, desired location. A robot may not move and settle in the same time step.
We wish to study swarms of autonomous robots whose goal is to attain uniform dispersion while minimizing three performance metrics: makespan, travel, and energy.
@@
Whereas travel measures the distance a robot has moved since its arrival, energy measures the amount of time during which it was active and consumed energy. Hence, a robot’s energy use continues increasing even when when it stays put, as long as it isn’t settled (e.g., a drone hovering in place):
@@
In context of this work, where all algorithms we consider have similar, neglibile, computational complexity, the primary energy expenditure differential comes from physical activities like movement, sensing, and maintaining basic operation—activities directly proportional to active time. This makes activity time a suitable proxy for energy consumption, an approach aligned with existing work in swarm robotics [22, 23].
These metrics are designed to capture both worst-case bounds on performance and average performance across the swarm. The maximal metrics ($T_{max}$, $E_{max}$) establish upper bounds on the resource requirements any robot may need, providing guarantees for hardware design and deployment. The total metrics ($T_{total}$, $E_{total}$) measure the swarm’s overall resource consumption, while makespan ($\mathcal{M}$) captures the collective completion time. These complementary perspectives—individual bounds and collective costs–are particularly important in swarm applications where both resource guarantees and overall efficiency matter.
@@
We are interested in studying autonomous robots with limited visibility that do not know their environment in advance. Primarily, we are interested in simple robots that have highly limited sensing, computation, and local signaling capabilities (as well as robots that cannot communicate at all). By “local signalling” we mean that each robot can broadcast a local visual or auditory signal by physical means which is received by all robots that sense it–see for example the light-based scheme in [46]. We assume recipients of a message know the relative position of the robot that sent the message. Other communication methods, such as radio communication, are also admissible as long as messages can be traced to the robot that originated them; but such schemes are much stronger than what our robots require. Let us parametrize the robots’ capabilities as follows:
@@
    Visibility range ($V$). A robot located at $p=(x,y)$ senses locations $q\in\mathbb{Z}^{2}$ at Manhattan distance $V$ or less from $p$. The robot knows the position relative to $p$ of every vertex $q\in\mathbb{Z}^{2}$ it senses (i.e., it knows $q-p$). We assume the robot can tell whether $q$ is $s$ and whether $q$ contains an obstacle (i.e., another robot or a wall). However, it cannot tell from sensing alone whether an obstacle is a wall, a settled robot, or an active robot.
@@
    Communication bandwidth ($B$) represents how many bits each robot can broadcast during the Move phase of each time step, after computing its next action. The broadcast is received by all robots that sense the broadcasting robot during their next Look phase. When $B=0$, robots cannot broadcast. For example, $B=1$ means a robot can broadcast a single bit (0 or 1) each time step, while $B=8$ allows broadcasting a byte of information.
@@
    Persistent state memory ($S$). Each robot has a state memory of $S$ bits that persists from time step to time step. At a given time $t$ each robot can be in any of $2^{S}$ states.
@@
In this work, we characterize a robot’s capabilities by a 3-tuple $(V,B,S)$. $(V,B,S)$ can be related to our robot’s hardware: a powerful platform might have large $V$, $B$, or $S$, whereas a simple robot has small $V$, $B$, and $S$. For example, $V$ is directly related to how powerful the robot’s perception hardware is (ranging from the 7 cm IR range on a Kilobot [46] to pocket-sized LiDARs that see 40 m); and basic perception plus a single on/off LED light gives $B=1$, while for higher $B$ we require a more sophisticated scheme can transmit more bits of information per time step.
The parameters $V$, $B$ and $S$ determine the possible algorithms robots can run. A central goal of ours is to study what algorithms become available to the robots as we increase the values $V$, $B$ or $S$. When can increasing these parameters improve the quality of the robots’ dispersion strategy, by enabling them to minimize a given performance metric?